suppose that a marksman hits the bull's eye 15,000 times outshots.If the next 4 shots are independent.Find the probability that

1.the next 4 shot hit the bull's eye.

2.Two of the next shot hits the bull's eye


1
Expert's answer
2020-07-20T12:37:03-0400

We assume that a marksman made n shots and 15000 of them were successful. Thus, a probability to make a successful outshot is "p=\frac{15000}{n}." Since we have four independent shots, we may use a binomial distribution with parameters n=4 and "p=\frac{15000}{n}" . Then, we receive:

  1. The probability that 4 outshots are successful is "(\frac{15000}{n})^4" .
  2. The probability that 2 of 4 shots are sucessful is "C_4^2(\frac{15000}{n})^2(1-\frac{15000}{n})^2=6(\frac{15000}{n})^2(1-\frac{15000}{n})^2" , where "C_4^2=\frac{4!}{2!2!}=6" is a binomial coefficent.

Answer:1. "(\frac{15000}{n})^4" 2. "6(\frac{15000}{n})^2(1-\frac{15000}{n})^2" ; "n" is a number of shots that a marksman made ( not including the last 4 shots).